# Multipartite Entanglement Distribution in Quantum Networks using Subgraph Complementations

### v4

@@ Abstract
Quantum networks are important for quantum communication and allow for several tasks such as quantum teleportation, quantum key distribution, quantum sensing, and quantum error correction. Graph states are a specific class of multipartite entangled states that can be represented by graphs. We propose a novel approach for distributing graph states across a quantum network. We show that the distribution of graph states can be characterized by a system of subgraph complementations, which we also relate to the minimum rank of the underlying graph and the degree of entanglement quantified by the Schmidt-rank of the quantum state. We analyze resource usage for our algorithm and show that it improves on the number of qubits, bits for classical communication, and EPR pairs utilized, as compared to prior work. In fact, the number of local operations and resource consumption for our approach scales linearly in the number of vertices. This produces a quadratic improvement in completion time for several classes of graph states represented by dense graphs, which translates into an exponential improvement by allowing parallelization of gate operations. This leads to improved fidelities in the presence of noisy operations, as we show through simulation in the presence of noisy operations. Common classes of graph states are classified along with their optimal distribution time using subgraph complementations. We find a close to optimal sequence of subgraph complementation operations to distribute an arbitrary graph state, and establish upper bounds on distribution time along with providing approximate greedy algorithms.
I Introduction
footnotetext: Main contactfootnotetext: Manning College of Information and Computer Sciences
I-A Motivation
@@
The theory of multipartite entanglement has been extensively used to study the distribution of graph states in a quantum network [1, 2, 3]. Creating entangled graph states efficiently is important for the development of quantum networks, having applications in areas such as quantum metrology [4], quantum error correction [5] and distributed quantum computing [6]. Graph states also form the basis of the one-way model of computation for quantum computers, which is universal and operates by performing single qubit measurements on an initially created resource graph state [7].
@@
In this paper, we first consider a quantum network in ideal conditions with no noisy communication and a limited number of perfect quantum memories. We address the problem of distributing a graph state under such conditions and present an algorithm to do so for a network having an arbitrary topology. We are in particular interested in gaining a better understanding of the limitations in resource consumption (of resources such as EPR pairs and number of qubits) and runtime associated with distributing graph states, and how they can be optimized. Secondly, we show that our algorithm leads to improved fidelities of the graph state in the presence of noisy operations.
I-B Related Work
@@
The distribution of arbitrary graph states in a quantum network has received attention recently. Some of the key parameters for distributing graph states are the number of Bell pairs consumed and the number of time steps required. Two approaches similar to our work (i.e. a noiseless network with an arbitrary topology) are provided by [2, 3]. The authors of [2] do not optimize resource consumption for an arbitrary graph state, but provide a framework for future work on this problem by analyzing resource consumption in a noiseless and lossless network. Their algorithm utilizes $(n-1)$ Bell pairs in one time step to distribute an $n$-qubit GHZ state, and $\frac{n(n-1)}{2}$ Bell pairs in $(n-1)$ time steps to distribute an arbitrary graph state with $n$ vertices [2]. Another approach using an identical setting [3] utilizes $\frac{3n^{2}-2n}{8}$ Bell pairs and takes $(n-1)$ time steps to distribute an arbitrary graph state. Both of these approaches assume that multiple gates can be applied on a single qubit in one time step. However, unlike these approaches, we consider the gate-based model of computation in which a qubit can only have one gate applied to it in one time step. We also assume that Bell pairs are distributed in advance between some central node and end nodes. This increases the distribution times of [2] and [3] by a factor of $n$, so they both require $O(n^{2})$ time steps.
@@
Other work [8, 9] considers additional factors such as the probabilistic nature of generating and the time for distributing Bell pairs, accounting for memory decoherence and noisy Bell state measurements. These factors significantly impact the rate and fidelity of distributing a graph state and are of interest to practically implement such protocols in a physical network.
@@
The algorithm provided by Avis et al. [8] (and Fischer, Towsley [3], under the assumption that the physical network has a star topology) improves on the work of [2] by proposing the following approach: 1) Prepare the desired $n$-qubit graph state locally at a central node; 2) distribute Bell pairs between the central node and each of the $n$ end nodes; 3) teleport each qubit in the central node to each of the end nodes, consuming all Bell pairs in the process.
@@
This approach utilizes $n$ Bell pairs, $O(|E|)=O(n^{2})$ time steps, and requires $2n$ qubits to be stored in the central node. We improve upon each one of these factors as described in the following section. Note that the distribution times for some of the above approaches can be improved to $O(n)$ by parallelizing gate operations. We also analyze the effect of incorporating parallelization in our approach and show a comparable speedup (within a $O(\log n)$ factor) for distributing all graph states, and an exponential improvement in distribution time, i.e. $O(\log n)$ time steps, for certain relevant classes of graphs.
I-C Summary of Contributions
@@
We propose a novel graph state distribution algorithm both with and without parallelization and analyze it first for the noiseless case. We then examine the effects of noise through simulations on various graph states. Our protocol is based on a specific operation on graph states, termed a subgraph complementation ($SC$), which should not be confused with local complementation. We then show that any graph state can be distributed in the network through a sequence of $SC$ operations. Along with introducing this protocol, we make the following contributions, improving upon prior work.
@@
  • We show that our graph state distribution algorithm utilizes $(n-1)$ Bell pairs and requires at most $(n-1)$ qubits at the central node. The number of Bell pairs is clearly optimal and the number of qubits at the central node approximately scales with the average degree of the vertices in the graph.
@@
  • We show that our algorithm requires $O(n)$ time steps and $O(n)$ bits of classical communication to distribute sparse graphs, as well as several relevant classes of dense graphs such as complete graphs (GHZ states) and complete bipartite graphs, but may take up to $O(n^{2})$ resources for some dense graphs.
@@
  • By parallelizing gate operations and using auxiliary qubits, we show that the distribution time improves to $O(\log n)$ time steps for the same classes of dense graphs. We obtain a general upper bound of $O(n\log n)$ time steps for arbitrary dense graphs and $O(n)$ for sparse graphs, as well as $O(n)$ time steps for random graphs (sparse and dense) with high probability.
  • We establish a relation between the number of $SC$ operations needed to distribute a graph state |$\ket{G}$ and the entanglement of |$\ket{G}$ (quantified by its average Schmidt-rank over all possible bipartitions).
@@
II Background
Graph states are a class of multipartite entangled states that can be represented by graphs, with each vertex representing a qubit and each edge representing a $CZ$ (controlled-$Z$) gate. A graph state $\left|G\right\rangle$ corresponding to the undirected graph $G=(V,E)$ is defined as
@@ II Background
[…]
@@
III Subgraph complementation protocol
We describe the protocol for the setting of no noise and deterministic Bell pair generation. This section will be followed by a discussion of $SC$ operations as part of our algorithm to distribute an arbitrary graph state.
III-A Single subgraph complementation
We first present a protocol for performing a single $SC$ operation on any subset of nodes $V_{\alpha}\subseteq V$ of the network, requiring $O(|V_{\alpha}|)$ time steps. This will serve as a building block for our graph state distribution algorithm.
The protocol by Avis et al. [8] distributes a state by locally creating it and then teleporting it, which is costly in memory usage. Roughly speaking, we get around this by constructing the desired graph state step by step through $SC$ operations which consist of local CZ gates (acting only on $n$ qubits as opposed to the $2n$ qubits in Avis et al. [8]) and local complementations.
More formally, consider a network with a central node $A$ that contains $k-1$ qubits $\{a_{i}:0\leq i<k-1\}$. The network also contains a set of $k-1$ nodes containing one qubit each $\{c_{i}:0\leq i<k-1\}$. Our setup assumes that $(k-1)$ Bell pairs of the form $(a_{i},c_{i}),0\leq i<(k-1)$, are already distributed between the nodes. Note that this defines two network graphs corresponding to the nodes (the physical graph) and qubits (the graph state) respectively.
The following protocol describes how to distribute a $k-clique$, or a fully connected graph of $k$ vertices, among the qubits in the set $\{a_{i}\}\bigcup\{c_{i}\}$ where qubits $a_{i}$ are in the central node $A$ and qubits $c_{i}$ are in the other end nodes. An example operation for $k=4$ can be found in Fig. 1.
Figure 1: 4-clique addition operation. If a 4-clique is the final state to be distributed, qubits $a_{1}$ and $a_{2}$ need to be measured out in the $Z$-basis at the end. Else, preserve the qubits for further $SC$ operations.
A
Subgraph complementation protocol on $V_{\alpha}$
@@ III-A Single subgraph complementation
    $LC(a_{0})$ — The resulting state corresponds to state (iv). This forms the complete graph for $V_{\alpha}$.
If a $k$-clique (corresponding to a $k$-qubit GHZ state) is the final state we want to distribute, then some qubits may need to be measured out in the $Z$-basis following the protocol (e.g. $a_{1}$ and $a_{2}$ in Fig. 1). However, if we want to distribute an arbitrary graph state, then it is beneficial to actually preserve connections of the form $(a_{i},c_{i})$ between qubits for carrying out additional $SC$ operations, in order to avoid sharing a new set of Bell pairs each time. This is described in more detail below.
III-B Sequence of subgraph complementations
In order to carry out another $SC$ operation after obtaining a state similar to state (iv) in Fig. 1, we need to restore direct connections $(a_{i},c_{i})$ between the central node and each of the end nodes, similar to the connections in state (i) in Fig. 1. We refer to this as an edge reset operation. This simply involves carrying out steps 1–3 in the subgraph complementation protocol in reverse order. This is illustrated in Fig. 2. The only difference between the new state (v) and state (i) is that the end nodes form a clique. The $SC$ protocol can then be carried out on a new set of vertices to add another clique, but since $CZ^{2}=I$, the edges of the clique get added modulo 2 (every two additions of the same edge removes it). More specifically, applying the sequence of operations 1–4 on any set of vertices $V_{\alpha}\subseteq V$ corresponds to performing another $SC$ operation on $V_{\alpha}$. This can be used to successively perform several $SC$ operations.
Finally, after performing some sequence of $SC$ operations, we would like to select which connections (edges) to keep or not to keep between the central node $A$ and the other nodes. Since $A$ is also part of the network, simply perform step 1 followed by step 2 of the $SC$ protocol to add any edges between $A$ (resp. $a_{i}$) and $C_{i}$ ($c_{i}$), and measure out all the other qubits.
@@ III-B Sequence of subgraph complementations
III-C Analysis
In order to complement a subgraph of $k$ vertices in a network, this protocol utilizes $k-1$ Bell pairs and $k-1$ qubits in the central node. A total of $(2k-2)$ $CZ$ operations and $k$ $LC$ operations are required by the basic protocol. This will be followed by a linear number of $Z$-measurements (maximum $k$), $LCs$ and $CZs$ (maximum $k-1$ in conjunction with each other). Thus, the number of time steps required to carry out this protocol is $O(k)$.
IV Graph State Distribution Algorithm
@@
IV-A Subgraph complementation system
@@
IV-B Time for subgraph complementation
For our analysis, we assume that Bell pairs are distributed in advance. We consider first the gate-based model of computation with perfectly noiseless quantum channels for communication (see Section VII for a numerical analysis in the presence of noise). Let $T(G,\mathcal{SC}_{sys})$ denote the time required by our protocol to distribute a graph state $\left|G\right\rangle$ according to subgraph complementation system $\mathcal{SC}_{sys}$. Denote by $t(k)$ the time required for a single $SC$ operation on $k$ vertices, $k\in\mathbb{N}$. It follows from the prior analysis section that $t(k)=\gamma k=O(k),$ for some $\gamma\in\mathbb{R^{+}}$, since time is proportional to the number of $CZ$ gates performed. This directly leads to an expression for $T(G,\mathcal{SC}_{sys})$ as stated below.
Theorem IV.1.

### v5

@@ Abstract
Quantum networks are important for quantum communication, enabling tasks such as quantum teleportation, quantum key distribution, quantum sensing, and quantum error correction, often utilizing graph states—a specific class of multipartite entangled states that can be represented by graphs. We propose a novel approach for distributing graph states across a quantum network. We show that the distribution of graph states can be characterized by a system of subgraph complementations, which we also relate to the minimum rank of the underlying graph and the degree of entanglement quantified by the Schmidt-rank of the quantum state. We analyze resource usage for our algorithm and show that it improves on the number of qubits, bits for classical communication, and EPR pairs utilized, as compared to prior work. In fact, the number of local operations and resource consumption for our approach scales linearly in the number of vertices. This produces a quadratic improvement in completion time for several classes of graph states represented by dense graphs, which translates into an exponential improvement by allowing parallelization of gate operations. This leads to improved fidelities in the presence of noisy operations, as we show through simulation in the presence of noisy operations. We classify common classes of graph states, along with their optimal distribution time using subgraph complementations. We find a sequence of subgraph complementation operations to distribute an arbitrary graph state which we conjecture is close to the optimal sequence [5], and establish upper bounds on distribution time along with providing approximate greedy algorithms.
1 Introduction
@@
The distribution of graph states in quantum networks has been extensively studied [14, 16, 12]. Creating entangled graph states efficiently is important for the development of quantum networks, having applications in areas such as quantum metrology [21], quantum error correction [15] and distributed quantum computing [8]. Graph states also form the basis of the one-way model of computation for quantum computers, which is universal and operates by performing single qubit measurements on an initially created resource graph state [19].
@@
In this paper, we first consider a quantum network in ideal conditions with no noisy communication and a limited number of perfect quantum memories. We address the problem of distributing a graph state under such conditions and present an algorithm to do so for a network having an arbitrary topology. We are particularly interested in gaining a better understanding of the limitations in resource consumption (of resources such as EPR pairs and number of qubits) and runtime associated with distributing graph states, and how they can be optimized. Secondly, we show that our algorithm leads to improved fidelities of the graph state in the presence of noisy operations.
1.1 Related Work
@@
The distribution of arbitrary graph states in a quantum network has received attention recently. Some of the key parameters for distributing graph states are the number of Bell pairs consumed and the number of time steps required. Two approaches similar to our work (i.e. a noiseless network with an arbitrary topology) are provided by [16, 12]. The authors of [16] do not optimize resource consumption for an arbitrary graph state, but provide a framework for future work on this problem by analyzing resource consumption in a noiseless and lossless network. Their algorithm utilizes (n1)(n-1) Bell pairs in one time step to distribute an $n$-qubit GHZ state, and $\frac{n(n-1)}{2}$ Bell pairs in (n1)(n-1) time steps to distribute an arbitrary graph state with $n$ vertices [16]. Another approach using an identical setting [12] utilizes $\frac{3n^{2}-2n}{8}$ Bell pairs and takes (n1)(n-1) time steps to distribute an arbitrary graph state. Both of these approaches assume that multiple gates can be applied on a single qubit in one time step, since this is negligible in comparison with the time for probabilistic Bell pair generation. However, unlike these approaches, we consider the gate-based model of computation in which a qubit can only have one gate applied to it in one time step. We also assume that Bell pairs are distributed in advance between some central node and end nodes. This increases the distribution times of [16] and [12] by a factor of $n$, so they both require $O(n^{2})$ time steps.
@@
Other work [2, 10, 6] considers additional factors such as the probabilistic nature of generating and the time for distributing Bell pairs, accounting for memory decoherence and noisy Bell state measurements. These factors significantly impact the rate and fidelity of distributing a graph state and are of interest to practically implement such algorithms in a physical network.
@@
The algorithm provided by Avis et al. [2], Cuquet, Calsamiglia [10], and Fischer, Towsley [12], (under the assumption that the physical network has a star topology) improves on the work of [16] by proposing the following approach, which we refer to as the Factory Node + Teleportation Algorithm or the FNT algorithm in the rest of the paper.
FNT graph state distribution algorithm
  1. 1.
    Prepare the desired $n$-qubit graph state locally at a central node
  2. 2.
    Distribute Bell pairs between the central node and each of the $n$ end nodes
  3. 3.
    Teleport each qubit in the central node to each of the end nodes, consuming all Bell pairs in the process
This algorithm is also described by Cuquet, Calsamiglia as the Bipartite A protocol [10]; they additionally provide an improved algorithm called Bipartite B which uses only $n$ qubits in the central node and fewer gates. However, they do not consider noise from local correction unitaries, which differs from our noise model as described in section 7. We show improved experimental results for fidelity in comparison to both models. These approaches utilize $n$ Bell pairs, $O(|E|)=O(n^{2})$ time steps, and require $2n$ or $n$ qubits to be stored in the central node. We improve upon each one of these factors as described in the following section.
@@
Note that the distribution times for some of the above approaches can be improved to $O(n)$ by parallelizing gate operations. We also analyze the effect of incorporating parallelization in our approach and show a comparable speedup for distributing all graph states, and an exponential improvement in distribution time, i.e. $O(\log n)$ time steps, for certain relevant classes of graphs.
1.2 Summary of Contributions
@@
We propose a novel graph state distribution algorithm both with and without parallelization and analyze it first for the noiseless case. We then examine the effects of noise through simulations on various graph states. Our algorithm is based on a specific operation on graph states, termed a subgraph complementation (SCSC)[5], which should not be confused with local complementation. We then show that any graph state can be distributed in the network through a sequence of $SC$ operations. Along with introducing this algorithm, we make the following contributions, improving upon prior work.
@@
  • We show that our graph state distribution algorithm utilizes (n1)(n-1) Bell pairs and requires at most (n1)(n-1) qubits at the central node. The number of Bell pairs is clearly optimal and the number of qubits at the central node approximately scales with the average degree of the vertices in the graph.
@@
  • We show that our algorithm requires $O(n)$ time steps and $O(n)$ bits of classical communication to distribute sparse graphs, as well as several relevant classes of dense graphs such as complete graphs (GHZ states) and complete bipartite graphs, but may take $O(n^{2})$ resources for some dense graphs.
@@
  • By parallelizing gate operations and using auxiliary qubits, we show that the distribution time improves to $O(\log n)$ time steps for the same classes of dense graphs. We obtain a general upper bound of $O(n\log n)$ time steps for arbitrary dense graphs and $O(n)$ for sparse graphs.
[…]
@@
2 Background
Graph states are a class of multipartite entangled states that can be represented by graphs, with each vertex representing a qubit and each edge representing a CZCZ (controlled-$Z$) gate. A graph state $\left|G\right\rangle$ corresponding to the undirected graph $G=(V,E)$ is defined as
@@ 2 Background
where $CZ_{ab}\left|+\right\rangle$ denotes an edge between two vertices $a,b\in V$, each of them initially in the $\left|+\right\rangle$ state, with $n=|V|$. Applying a $CZ$ gate twice removes an edge, i.e. $CZ^{2}=I$. Note that $CZ$ gates commute with each other and are also symmetric, so edges are undirected.
@@
3 Subgraph complementation operation
We describe the subgraph complementation operation for the setting of no noise and deterministic Bell pair generation. This section will be followed by a discussion of $SC$ operations as part of our algorithm to distribute an arbitrary graph state.
3.1 Single subgraph complementation
We first present an algorithm for performing a single $SC$ operation on any subset of nodes $V^{\prime}\subseteq V$ of the network, requiring $O(|V^{\prime}|)$ time steps. This will serve as a building block for our graph state distribution algorithm.
Roughly speaking, we construct the desired graph state step by step through $SC$ operations that consist of local CZ gates (acting only on n1n-1 qubits as opposed to the $2n,n$ qubits in [2, 10]) and local complementations.
More formally, consider a network with a central node $A$ that contains k1k-1 qubits $\{a_{i}:0\leq i<k-1\}$. The network also contains a set of k1k-1 nodes $\{C_{i}\}_{i=0}^{k-2}$, each containing one qubit $\{c_{i}\}_{i=0}^{k-2}$. The qubit $a_{0}$ will be used as a ‘central’ representative qubit for node $A$. Our setup assumes that (k1)(k-1) Bell pairs of the form $(a_{i},c_{i}),i=0,\ldots,k-2$, are already distributed between the nodes. Note that this defines two network graphs corresponding to the nodes (the physical graph) and qubits (the graph state) respectively. We make no assumptions about the underlying topology or connectivity of the network, since our algorithm only requires Bell pairs to be distributed between an arbitrary central node and all the other nodes. This can be achieved starting from any connected physical graph, though the optimal method to do so may be nontrivial to find. For the analysis in this paper we focus on a star network topology, or equivalently, we exclude all the resources and contributions to the noise required to connect the end users to the central node. We focus on this setting since a star network topology is considered to be a reasonable architecture that is implementable in the near term.
The simplest $k$-qubit graph state we can distribute using one $SC$ operation is the $k$-GHZ state or a complete graph state on $k$ qubits (i.e. a $k$-clique). The following algorithm describes how to distribute a $k$-clique among all the nodes $V=\{A\}\bigcup\{C_{i}\}_{0}^{k-2}$. This involves carrying out one $SC$ operation on k1k-1 nodes on $V^{\prime}=\{C_{i}\}$ followed by measurements in $A$. An example operation for $k=4$ can be found in Fig. 1.
Figure 1: 4-clique addition operation. The SC operation is first carried out on $V^{\prime}=\{C_{0},C_{1},C_{2}\}$ containing qubits $\{c_{0},c_{1},c_{2}\}$.
If a 4-clique is the final state to be distributed, qubits $a_{1}$ and $a_{2}$ need to be measured out in the $Z$-basis at the end. Else, preserve the qubits for further $SC$ operations. A
Subgraph complementation operation on $V^{\prime}$
@@ 3.1 Single subgraph complementation
    $LC(a_{0})$ — The resulting state corresponds to state (iv). This complements the induced graph on $V^{\prime}$, which was initially an empty graph.
If a $k$-clique (corresponding to a $k$-qubit GHZ state) is the final state we want to distribute, then some qubits may need to be measured out in the $Z$-basis at the end (e.g. $a_{1}$ and $a_{2}$ in Fig. 1). However, if we want to distribute an arbitrary graph state, then it is beneficial to actually preserve some or all connections of the form $(a_{i},c_{i})$ between qubits for carrying out additional $SC$ operations, in order to avoid sharing a new set of Bell pairs each time. This is described in more detail below.
3.2 Sequence of subgraph complementations
In order to carry out another $SC$ operation after obtaining a state similar to state (iv) in Fig. 1, we need to restore direct connections $(a_{i},c_{i})$ between the central node and each of the end nodes, similar to the connections in state (i) in Fig. 1. We refer to this as an edge reset operation. This simply involves carrying out steps 1–3 in the subgraph complementation operation in reverse order. This is illustrated in Fig. 2. The only difference between the new state (v) and state (i) is that the end nodes form a clique. The $SC$ operation can then be carried out on a new set of vertices to add another clique, but since $CZ^{2}=I$, the edges of the clique get added modulo 2 (every two additions of the same edge removes it). More specifically, applying the sequence of operations 1–4 on any set of vertices $V^{\prime}\subseteq V$ corresponds to performing another $SC$ operation on $V^{\prime}$. This can be used to successively perform several $SC$ operations.
Finally, after performing some sequence of $SC$ operations, we would like to select which connections (edges) to keep or not to keep between the central node $A$ and the other nodes. Since $A$ is also part of the network, simply perform step 1 followed by step 2 of the $SC$ operation to add any edges between $A$ (resp. $a_{i}$) and $C_{i}$ ($c_{i}$), and measure out all the other qubits.
@@ 3.2 Sequence of subgraph complementations
3.3 Analysis
In order to carry out an $SC$ operation on $k^{\prime}=k-1$ vertices in a network that includes the central node, we utilize $k^{\prime}$ Bell pairs and $k^{\prime}$ qubits in the central node. A total of $(2k^{\prime}-2)$ CZCZ operations and $k^{\prime}$ LCLC operations are required by the SC operation 3.1. This is followed by up to $k^{\prime}$ $Z$-measurements, up to $k^{\prime}-1$ LCsLCs and $CZs$ in conjunction with each other depending on the number of edge resets. Thus, the number of time steps required to carry out a subgraph complementation followed by an edge reset is bounded by $6k^{\prime}-4=O(k^{\prime})$.
4 Graph State Distribution Algorithm
@@
4.1 Subgraph complementation system
@@
[…]
Graph state distribution algorithm via $\mathcal{SC}_{sys}$
  1. 1.
    Initialization. Given a fixed $\mathcal{SC}_{sys}$ for $G$, start with a fixed set of n1n-1 Bell pairs between an arbitrary central node and all the other nodes in the network.
  2. 2.
    Apply subgraph complementation operations. For each $V_{i}\in\mathcal{SC}_{sys}$, perform the corresponding $SC$ operation on $V_{i}$, followed by an edge reset on $V_{i}$. Each operation complements a $|V_{i}|$-size clique for the vertices in $V_{i}$ in $G$. The order of the $|\mathcal{SC}_{sys}|$ operations does not affect the final graph state.
  3. 3.
    Last edge reset. After the last SCSC operation, add or remove the appropriate edges between the central node and other nodes in $G$ (the first 3 steps of the $SC$ operation 3.1 allows us to add any set of edges $(a_{0},c_{i})$ as in Fig. 1, and we can measure out qubit $a_{i}$ to remove the edge $(a_{0},c_{i})$).
  4. 4.
[…]
4.2 Time for subgraph complementation
For our analysis, we assume that Bell pairs are distributed in advance. We consider first the gate-based model of computation with perfectly noiseless quantum channels for communication (see Section 7 for a numerical analysis in the presence of noise). Let $T(G,\mathcal{SC}_{sys})$ denote the time required by our algorithm to distribute a graph state $\left|G\right\rangle$ according to subgraph complementation system $\mathcal{SC}_{sys}$. Denote by $t(k)$ the time required for a single $SC$ operation on $k$ vertices, $k\in\mathbb{N}$. It follows from the prior analysis section that $t(k)=\gamma k=O(k),$ for some $\gamma\in\mathbb{R^{+}}$, since time is proportional to the number of $CZ$ gates performed. This directly leads to an expression for $T(G,\mathcal{SC}_{sys})$ as stated below.
Theorem 4.1.